### v2

# Phase space noncommutativity, power-law inflation and quantum cosmology

Abstract
Considering an arbitrary dimensional FLRW universe in the framework of a generalized Sáez–Ballester (SB) theory, we establish a noncommutative (NC) cosmological model. We concentrate on the predictions of NC model and compare them with their commutative counterparts in both the classical and quantum regimes. For the classic case, taking a very small NC parameter, we apply two different methods to analyze the model features. First, we show through numerical analysis that our NC model is a successful inflationary model capable of overcoming the graceful exit and horizon problems. Furthermore, the NC traces are visible the late time, which supports the UV/IR mixing characteristic of the NC models. In the second method, we show that our NC model can correspond to the previously developed NC inflationary models. In the commutative quantum case, we obtain an exact wave function and then use the WKB approximation to show that the solutions of the corresponding classical regime are recovered. Finally, with regard to the NC quantum level, we focus on the special case for which we show that a constant of motion exists. The latter helps us to conveniently transform the corresponding complicated NC-WDW equation into an ordinary differential equation, which can be easily solved numerically for the general case or analytically for some special cases. The resultant solutions show a damping behavior in the wave function associated with the proposed NC model, which may be important in determining the viable initial states for the very early universe.
keywords:
Sáez–Ballester theory , Noncommutativity , Inflation , Quantum cosmology , WDW equation , FLRW Cosmology , UV/IR mixing , WKB approximation
1 Introduction
Models based on noncommutative (NC) theories are still of interest. The main motivation for employing noncommutativity of the canonical type may be traced back to the investigations that establish its relationship with string and M theory [1]. In addition, the works in semi–classical gravity [2] could be related to NC field theories. Furthermore, the application of NC theories allows for the consideration of several fascinating aspects, such as physics at extremely small distances [3], non-locality, IR/UV mixing [4], Lorentz violation [5], the NC Landau problem [6] and the quantum Hall effect [7].
Many NC scenarios have been established as cosmological applications, see for example Newtonian cosmology [8], quantum cosmology [9], perturbation cosmology and NC inflation cosmology [10]. Let us be more precise. NC cosmological models have the ability to solve some of the most important outstanding problems in cosmology by opening up new pathways for studying the fundamental nature of the cosmos, see, for instance [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] and references therein. Modified dispersion relations, NC inflation, dark matter and dark energy, the cosmic microwave background, and quantum gravity effects are just a few of the relevant issues that the NC cosmological scenarios can predict appropriately, see for instance [22, 23] and references therein.
Quantum effects are expected to play a key role in the cosmic evolution of the universe. Unifying GR with quantum mechanics has been a outstanding problem, for which NC geometry could provide insights and aid in the development of a full theory of quantum gravity. Among the various candidates for quantum gravity, many models of NC gravity have been proposed [10, 9]. Most NC models assume a homogeneous universe where the metric (as well as the applied scalar field) depends only on time. The resulted model is a finite-dimensional configuration space called a minisuperspace.
The main objective of this work is to investigate of the traces of noncommutativity in early times in (semi-) classical and quantum regimes. Concretely, we propose a dynamical noncommutativity for a cosmological model in which we consider a deformation of minisuperspace rather than a deformation of spacetime. Let us be more precise. In the classical regime (for both the commutative and the NC models), as well as in the commutative quantum scenario, we will consider a specific potential, so an extended version of the standard power-law inflationary model provided in Ref. [24] is elaborated. Subsequently, in the NC classical scenario, we will demonstrate not only that our model has the characteristics of a successful inflationary model, but also, using another method for solving differential equations, that our model corresponds to an NC inflationary model established in Ref. [25]. Finally, to explore solutions for our quantum NC model, we focus on a specific case where the potential takes a constant value and the chosen dynamic noncommutativity is assumed to be the simplest case. To present our analysis for this case we use Seiberg-Witten (SW) map, which helps convert the NC system into an appropriately modified commutative one.
The paper is organized as follows. In the next section, considering the FLRW metric, we proposed an NC cosmological model in the context of the Sáez–Ballester (SB) theory in arbitrary dimensions. In sections 3 and 4, we study the commutative and NC SB cosmology in classical regime and compare the corresponding results. In sections 5 and 6, we investigate the quantum version of the models mentioned. Due to the complexity of the field equations for the NC quantum model, we restrict ourselves to a particular case. In Section 7, we provide an overview of the models presented and highlight some important findings.
2 Noncommutative SB cosmology
Let us consider the $D$-dimensional Friedmann–Lemaître–Robertson–Walker (FLRW) metric as the background geometry:
$ds^{2}=-N(t)dt^{2}+a^{2}(t)\left(\frac{dr^{2}}{1-\mathcal{K}r^{2}}+r^{2}d\Omega_{{}_{D-2}}^{2}\right),$ (1)
where $\mathcal{K}=-1,0,1$; $N(t)$ and $a(t)$ denote the lapse function and the scale factor, respectively; $d\Omega_{{}_{D-2}}^{2}=d\theta_{1}^{2}+sin^{2}\theta_{1}d\theta_{2}^{2}+...+sin^{2}\theta_{1}...sin^{2}\theta_{D-3}d\theta_{D-2}^{2}$ for $D\geq 3$.
Moreover, instead of the conventional Sáez–Ballester (SB) theory [39], we consider a generalized version of it, which contains a scalar potential, in $D$ dimensions as an underlying gravitational framework:
${\cal S}_{{}_{\rm SB}}^{{}^{(D)}}=\int d^{{}^{\,D}}\!x{\cal L}=\int{\rm d}td^{{}^{D-1}}\!x\sqrt{-g}\,\Big{[}R^{{}^{(D)}}-{\cal W}\phi^{n}\,g^{\alpha\beta}\,({\nabla}_{\alpha}\phi)({\nabla}_{\beta}\phi)-V(\phi)\,-2\chi\varrho(a)\Big{]},$ (2)
where ${\cal W}$ and $n$ are the parameters of the model; $\chi\equiv 8\pi G$; the homogeneous scalar filed $\phi$ is minimally coupled to the Ricci scalar; the Greek indices run from zero to $D-1$; the covariant derivative is denoted by $\nabla$; $g$ and $R^{{}^{(D)}}$ are the determinant and Ricci scalar associated with the $D$-dimensional metric $g_{\alpha\beta}$, respectively and we used the units where $c=1=\hbar$. From now on, we will consider only the spatially flat FLRW universe.
The Ricci scalar corresponding to the spatially flat FLRW metric (1) is given by
$R^{{}^{(D)}}=\frac{2(D-1)}{N^{2}}\left[\frac{\ddot{a}}{a}+\frac{D-2}{2}\left(\frac{\dot{a}}{a}\right)^{2}-\left(\frac{\dot{N}}{N}\right)\left(\frac{\dot{a}}{a}\right)\right].$ (3)
Substituting $R^{{}^{(D)}}$ from (3) into (2), we obtain
${\cal L}=-Na^{D-1}\left[\frac{(D-1)(D-2)}{N^{2}}\left(\frac{\dot{a}}{a}\right)^{2}-\frac{{\cal W}\phi^{n}}{N^{2}}\dot{\phi}^{2}+V(\phi)+2\chi\varrho(a)\right],$ (4)
where $\dot{A}\equiv dA/dt$.
The conjugate momenta associated with the scale factor and scalar field are:
$\Pi_{a}$ $=$ $-2(D-1)(D-2){N}^{-1}a^{D-3}\dot{a},$ (5)
$\Pi_{\phi}$ $=$ $2{\cal W}{N}^{-1}a^{D-1}\phi^{n}\dot{\phi}.$ (6)
From using the standard relationship between the Lagrangian and the classical Hamiltonian $\mathcal{H}$, it is straightforward to show that
$\mathcal{H}=-\frac{{\Pi^{2}_{a}}}{4(D-1)(D-2)a^{D-3}}+\frac{\Pi^{2}_{\phi}}{4{\cal W}a^{D-1}\phi^{n}}+a^{D-1}\left[V(\phi)+2\chi\varrho(a)\right],$ (7)
where, for the sake of technical simplicity, we used the particular gauge $N=1$, which is only relevant for the classic level. It should be noted that, at the quantum level, the lapse function does not appear in the field equations.
Let us propose a NC cosmological model considering a deformed Poisson algebra as
$\{a,\Pi_{a}\}=1=\{\phi,\Pi_{\phi}\},\hskip 14.22636pt\{\Pi_{a},\Pi_{\phi}\}=\theta\phi^{\kappa},$ (8)
where the other Poisson brackets vanish, $\theta={\rm constant}$ is the NC parameter and
$\kappa\equiv\frac{D(n+2)}{D-2}-1,$ (9)
which is proposed based on the dimensional analysis; see also Refs. [11, 26, 27] for more information on why it is interesting to propose noncommutativity between momenta.
Employing relations (8) and
$\left\{\Pi_{a},f(\Pi_{a},{\Pi}_{\phi})\right\}=\theta\phi^{\kappa}\frac{\partial f}{\partial{\Pi}_{\phi}},\hskip 22.76219pt\left\{{\Pi}_{\phi},f(\Pi_{a},{\Pi}_{\phi})\right\}=-\theta\phi^{\kappa}\frac{\partial f}{\partial\Pi_{a}},$ (10)
where $f$ is an arbitrary function of the momenta. The classical equations of motion for the phase space variables $a$, $\phi$, $\Pi_{a}$ and $\Pi_{\phi}$ are given by
$\dot{a}$ $=$ $-\frac{1}{2(D-1)(D-2)}a^{3-D}\Pi_{a},$ (11)
$\dot{\Pi}_{a}\!\!\!$ $=$ $-\frac{D-3}{4(D-1)(D-2)}a^{2-D}\Pi_{a}^{2}+\frac{D-1}{4{\cal W}}a^{-D}\phi^{-n}\Pi_{\phi}^{2}$
-- $(D-1)a^{D-2}\left[V(\phi)+2\chi\varrho(a)\right]-2\chi a^{D-1}\frac{d\varrho(a)}{da}+\frac{\theta}{2{\cal W}}a^{1-D}\phi^{\kappa-n}\Pi_{\phi},$
$\dot{\phi}\!\!\!$ $=$ $\frac{1}{2{\cal W}}a^{1-D}\phi^{-n}\Pi_{\phi},$ (13)
$\dot{\Pi}_{\phi}$ $=$ $\frac{n}{4{\cal W}}a^{1-D}{\phi}^{-(n+1)}\Pi_{\phi}^{2}-a^{D-1}\frac{dV(\phi)}{d\phi}+\frac{\theta}{2(D-1)(D-2)}a^{3-D}\phi^{\kappa}\Pi_{a}$
and
$\mathcal{H}=-\frac{{a^{3-D}\Pi^{2}_{a}}}{4(D-1)(D-2)}+\frac{a^{1-D}\phi^{-n}\Pi^{2}_{\phi}}{4{\cal W}}+a^{D-1}\left[V(\phi)+2\chi\varrho(a)\right]\approx 0.$ (15)
After some manipulations, we obtain the classical equations of motion:
$(D-1)H^{2}$ $=$ $\rho_{{}_{\phi}}+\frac{2\chi}{(D-2)}\varrho(a),$ (16)
$2\frac{\ddot{a}}{a}+(D-3)H^{2}\!$ $=$ $\!-\left(p_{{}_{\phi}}+p_{{}_{\rm nc}}\right)+\frac{2\chi}{D-2}\left[\varrho(a)+\frac{a}{D-1}\frac{d\varrho(a)}{da}\right],$ (17)
$\ddot{\phi}+(D-1)H\dot{\phi}\!$ $+$ $\!\frac{n}{2}{\phi}^{-1}\dot{\phi}^{2}+\frac{1}{2{\cal W}}{\phi}^{-n}\frac{dV(\phi)}{d\phi}+\frac{\theta a^{2-D}H\phi^{\kappa-n}}{2{\cal W}}=0,$ (18)
where $H\equiv\dot{a}/a$; $\rho_{{}_{\phi}}$ and $p_{{}_{\phi}}$ denote the energy density and pressure of the homogeneous scalar field:
$\rho_{{}_{\phi}}$$\equiv$ $\frac{1}{D-2}\left[{\cal W}{\phi}^{n}\dot{\phi}^{2}+V(\phi)\right],$ (19)
$p_{{}_{\phi}}$$\equiv$ $\frac{1}{D-2}\left[{\cal W}{\phi}^{n}\dot{\phi}^{2}-V(\phi)\right].$ (20)
Moreover, the quantity $p_{{}_{\rm nc}}$, which refers to an effective pressure term arising due to the presence of the chosen dynamical deformation (8), was defined as
$p_{{}_{\rm nc}}\equiv\frac{\theta}{(D-1)(D-2)}a^{2-D}\phi^{\kappa}\dot{\phi}.$ (21)
Obtaining analytic solutions for the NC equations (16)-(18) seems very complicated, so in Section 4 we will present predictions for some physical quantities using numerical analysis. In addition, solving the NC-WDW equation for the corresponding quantum cosmological model is also not easy. Therefore, in Section 6, we will only focus on a particular case.
In this work we are interested in studying the early universe where only the scalar field dominates the dynamics. Therefore, from now on we restrict ourselves to cosmological models that do not take ordinary matter into account. In this case, it is easy to show that the following conservation equation is identically satisfied:
$\Delta\equiv\dot{\rho}_{{}_{\phi}}+(D-1)H\left(\rho_{{}_{\phi}}+p_{{}_{\phi}}+p_{{}_{\rm nc}}\right)=0.$ (22)
Among the differential equations (16)-(18) and (22), only two of them are independent. To specify these two, we would like to focus on the following equations:
$H=\epsilon\sqrt{\frac{\rho_{{}_{\phi}}}{D-1}},$ (23)
$\frac{\dot{\rho}_{{}_{\phi}}}{\sqrt{\rho_{{}_{\phi}}}}=-\frac{\sqrt{D-1}}{D-2}\left[2{\cal W}{\phi}^{n}\dot{\phi}^{2}+\frac{\theta}{D-1}a^{2-D}\phi^{\kappa}\dot{\phi}\right],$ (24)
where $\epsilon\equiv\pm 1$.
Let us assume 111 The justifications for choosing this type of potential are outlined in Ref. [28]. Except for Section 6, this potential will be used in all sections of this paper.
$V(\phi)=\left(\frac{2}{\Gamma}-{\cal W}\right)\phi^{n}\dot{\phi}^{2},$ (25)
where $\Gamma$ is a dimensionless constant.
By substituting $V(\phi)$ from (25) into equations (19), (23) and (24), we obtain
$\rho_{{}_{\phi}}=$ $\>\frac{2{\phi}^{n}\dot{\phi}^{2}}{(D-2)\Gamma},$ (26)
$H=$ $\>\epsilon h_{0}\phi^{\frac{n}{2}}\dot{\phi},$ (27)
$\frac{d}{dt}\left(\phi^{n}\dot{\phi}^{2}\right)=$ $\>-\epsilon\sqrt{\frac{(D-1)\Gamma}{2(D-2)}}\left[2{\cal W}{\phi}^{n}\dot{\phi}^{2}+\frac{\theta}{(D-1)}a^{2-D}\phi^{\kappa}\dot{\phi}\right]\phi^{\frac{n}{2}}\dot{\phi},$ (28)
where we have used
$\sqrt{\phi^{n}\dot{\phi}^{2}}=\epsilon\phi^{\frac{n}{2}}\dot{\phi},\hskip 42.67912pth_{0}\equiv\sqrt{\frac{2}{(D-1)(D-2)\Gamma}}.$ (29)
Equations (27) and (28) are the independent equations associated with our proposed NC model for the chosen potential (25).
3 Commutative classical cosmology
In this case, substituting $\theta=0$ into equation (28), we obtain
$\left(\frac{1}{\phi^{\frac{3n}{2}}\dot{\phi}^{3}}\right)\frac{d}{dt}\left(\phi^{n}\dot{\phi}^{2}\right)=\epsilon\zeta,$ (30)
where
$\zeta\equiv-{\cal W}\sqrt{2\Gamma\left(\frac{D-1}{D-2}\right)}=-(D-1){\cal W}\Gamma h_{0}.$ (31)
We can easily obtain a solution for (30) as
$\phi^{\frac{n}{2}}\dot{\phi}=-\left(\frac{2\epsilon}{\zeta}\right)\frac{1}{t},$ (32)
where the integration constant has been set equal to zero.
Equation (32) gives
[…]
Substituting $\phi^{\frac{n}{2}}\dot{\phi}$ from (32) into equations (25), (26) and (27), we can easily obtain the potential, energy density and the scale factor in terms of the cosmic time:
$V(t)$ $=$ $V_{i}\left(\frac{t}{t_{i}}\right)^{-2},\hskip 51.21495ptV_{i}\equiv\left[\frac{4(2-{\cal W}\Gamma)}{\zeta^{2}\Gamma}\right]\frac{1}{t_{i}^{2}},$ (34)
$\rho_{\phi}(t)$ $=$ $\left[\frac{8}{(D-2)\Gamma\zeta^{2}}\right]\,\frac{1}{t^{2}},$ (35)
$a(t)$ $=$ $a_{i}\left(\frac{t}{t_{i}}\right)^{\beta},\hskip 51.21495pt\beta\equiv\frac{2\epsilon}{(D-1){\cal W}\Gamma},$ (36)
which are valid for all values of $n$. In (36), $a_{i}$ is an integration constant. For $\epsilon=1$ and $\beta>1$ one can obtain a power-law inflation [24, 28, 29] at early times. It should be noted that, because $\beta$ is a constant in this scenario, assuming that it takes values greater than one, the universe accelerates indefinitely. The number of e-folds, $\mathcal{N}$, during the period of inflation is:
[…]
which depends on the number of the dimensions of the universe and ${\cal W}$. Such a solution has been also obtained using other gravitational frameworks [30, 31, 32].
Employing relations (33) and (34), we can obtain the potential in terms of the scalar field:
$V(\phi)=\left\{\begin{array}[]{c}V_{i}{\exp}\left[\left(\frac{2\epsilon\zeta}{n+2}\right)\left(\phi^{\frac{n+2}{2}}-\phi_{i}^{\frac{n+2}{2}}\right)\right],\hskip 34.1433pt{\rm for}\hskip 28.45274ptn\neq-2,\\ \\ V_{i}\left(\frac{\phi}{\phi_{i}}\right)^{\epsilon\zeta},\hskip 108.12047pt{\rm for}\hskip 28.45274ptn=-2.\end{array}\right.$ (38)
Using relations (32) and (34), we showed that
$p_{\phi}(t)=\left[\frac{8({\cal W}\Gamma-1)}{\zeta^{2}\Gamma}\right]\,\frac{1}{t^{2}},\hskip 36.98857pt\forall n.$ (39)
In the next section we will analyze the behavior of some of the physical quantities obtained above and compare them with their corresponding NC counterparts.
4 Noncommutative classical SB cosmology
In this section we assume that the NC parameter is nonzero and takes very small values; clearly, when it vanishes, all field equations are reduced to those of the commutative case.
To specify the behavior of physical quantities for the classical NC model, we can consider two methods as follows.
4.1 First method: numerical investigation
In this method, we first solve the equation (27) to relate the scale factor to the scalar field. If we then plug the resulting relation into (28), we get a differential equation for just $\phi(t)$. It turns out that the solution of such a complicated equation is not analytically feasible. Therefore we have to use a numerical method to analyze the behavior of the required quantities. Because we will be studying the dynamics of the problem below, we will refrain from illustrating the behavior of all physical quantities and will only briefly describe the results here.
Our numerical investigations show that we can reach an inflationary epoch at earlier times, such that after a short time the scale factor transitions gracefully from an acceleration phase and enters to a deceleration phase, which is due to a radiation–dominated universe. In addition, the nominal condition is relevant for the causal physics of inflation is met. Other more appropriate methods, which are discussed below, can be utilised to demonstrate the above results.
To prove that our model is a successful inflationary model, it is easy to provide its corresponding dynamical system as
$\dfrac{d\tilde{a}}{da}=\left[1-\frac{\mathcal{W}}{(D-2)h_{0}^{2}}\right]\left(\frac{\tilde{a}}{a}\right)-\frac{\theta a^{2-D}\left[\phi_{i}^{\frac{n+2}{2}}+\frac{n+2}{2\alpha}\ln\left(\frac{a}{a_{i}}\right)\right]^{\frac{D+2}{D-2}}}{2(D-1)(D-2)h_{0}},\hskip 8.53581ptn\neq-2,$ (40)
and
$\dfrac{d\tilde{a}}{da}=\left[1-\frac{\mathcal{W}}{(D-2)h_{0}^{2}}\right]\left(\frac{\tilde{a}}{a}\right)-\frac{\epsilon\theta a^{2-D}}{2(D-1)(D-2)h_{0}},\hskip 48.36967ptn\neq-2,$ (41)
where $\tilde{a}=\dot{a}$, and we have used (17) and (27). Our numerical endeavors have shown that the model is very sensitive to the values of the parameters and the constants of integration ($a_{i}$ and $\phi_{i}$) and initial conditions (ICs). However, we always have an inflationary epoch that corresponds to certain sets of reasonable values. Let us focus more precisely on the phase portrait of equations (40) and (41), by which we compare the commutative model and the corresponding NC model using the same values for the common parameters, integration constants and ICs. As the left panels of figures 1 and 2 show, for the commutative case, $\dot{a}$ always decreases as scale factor increases, indicating a decelerating universe. For an NC model fitted to very small values of the NC parameter, we get an interesting behavior for the scale factor. Specifically, as shown in Figures 1 and 2 (see right panels), we get an inflationary epoch in very early times. Thereafter, there is an suitable mechanism for exiting from the accelerating phase gracefully and thus a transition into the deceleration phase. It is worth noting that the application of the exponential potential is useful in most classical models to obtain exact solutions for an inflationary epoch. However, these models, without dictating some additional conditions, usually do not have a suitable mechanism to end the inflationary phase. We emphasize that the problem has been solved in our NC model.
It is important to note that the mentioned statements are also valid for $n=-2$, see figure 2.
Figure 1: Phase space portrait associated with the commutative case (left panel) and NC case (the right panel assuming $\theta=-0.0001$), respectively, according to equation (40). We have assumed $n=3$, ${\cal W}=2$, $\Gamma=0.8$, $18\pi G=$ 1, $a_{i}=0.005$, $\phi_{i}=1$ and $D=4$.
Figure 2: Phase space portrait associated with the commutative case (left panel) and NC case (the right panel assuming $\theta=-0.01$) for the case $n=-2$, according to equation (41). We have assumed ${\cal W}=0.09$, $\Gamma=9.9$, $18\pi G=$ 1, and $D=4$.
Since we cannot reconstruct the Lagrangian belonging to our NC model, let us compare our model with the well-known Starobinsky model [33] (see also [34]) at the level of the associated field equations. Employing equations (17) and (27), it is easy to show that, the evolution of the scale factor for our NC model for $n\neq-2$ is
$\frac{\dddot{a}}{a}$ $+$ $\left[\frac{2{\cal W}-3(D-2)h_{0}^{2}}{(D-2)h_{0}^{2}}\right]\left(\frac{\dot{a}\ddot{a}}{a^{2}}\right)-2\left[\frac{{\cal W}-(D-2)h_{0}^{2}}{(D-2)h_{0}^{2}}\right]\left(\frac{\dot{a}}{a}\right)^{3}$ (42)
$+$ $\frac{\theta a^{2-D}}{2(D-1)(D-2)h_{0}}\left[\phi_{i}^{\frac{n+2}{2}}+\frac{n+2}{2\alpha}\ln\left(\frac{a}{a_{i}}\right)\right]^{\frac{D+2}{D-2}}\Bigg{\{}\frac{\ddot{a}}{a}+\left(\frac{\dot{a}}{a}\right)^{2}$
$\times$ $\left(\frac{(D+2)(n+2)}{2(D-2)h_{0}}\left[\phi_{i}^{\frac{n+2}{2}}+\frac{n+2}{2\alpha}\ln\left(\frac{a}{a_{i}}\right)\right]^{-1}-(D-1)\right)\Bigg{\}}=0$
and for $n=-2$ is
$\frac{\dddot{a}}{a}$ $+$ $\left[\frac{2{\cal W}-3(D-2)h_{0}^{2}}{(D-2)h_{0}^{2}}\right]\left(\frac{\dot{a}\ddot{a}}{a^{2}}\right)-2\left[\frac{{\cal W}-(D-2)h_{0}^{2}}{(D-2)h_{0}^{2}}\right]\left(\frac{\dot{a}}{a}\right)^{3}$ (43)
$+$ $\frac{\epsilon\theta a^{2-D}}{2(D-1)(D-2)h_{0}}\left[\frac{\ddot{a}}{a}+(1-D)\left(\frac{\dot{a}}{a}\right)^{2}\right]=0.$
For the special case that our model is reduced to the Einstein scalar field system with canonical kinetic term, the equations (42) and (43) can correspond to the evolution equation for the scale factor, which was obtained in Ref. [34] and justify the presence of the NC parameter and the integration constants (of our model) by the presence of the coefficients introduced in [34].
4.2 Second method
In this subsection, we would like to explore another interesting approximate solution for our NC model. Let us examine the intriguing NC inflationary model [25]:
$a(t)=a_{i}e^{\sqrt{\frac{\pi\Theta}{2}}H_{i}\text{erf}\left(\frac{t}{\sqrt{2\Theta}}\right)},$ (44)
where $H_{i}$ is the value of the Hubble parameter at $t=t_{i}$, $\Theta\propto\theta^{2}$, $a_{i}$ is a constant and
[…]
Substituting $a(t)$ from (44) into (27), we obtain
$\phi(t)=\left[\frac{n+2}{4}\left(C\pm\sqrt{2\pi\Theta}\,\text{erf}\left(\frac{t}{\sqrt{2\Theta}}\right)\right)\right]^{\frac{2}{n+2}},$ (46)
where $n\neq-2$ and $C$ is a constant of integration.
For such a solution, as an example, in figure 3, we have plotted the qualitative behavior of some relevant physical quantities in terms of time varying from $\infty$- ∞ to +$\infty$+ ∞. As seen, the Hubble parameter reaches its maximum at $t=0$, while $\ddot{a}$ changes sign and turn to be negative when the comoving Hubble length $(aH)^{-1}$ and the effective (NC) pressure $p_{{}_{\rm nc}}$ reach their minimum. Our numerical results reveal that the smaller $\Theta$, the better equation (22) (or (28)) is satisfied, see for instance figure 4.
Figure 3: Left panel: The time behavior of the scale factor $a$ (solid curve), $\ddot{a}$ (dashed curve) and the Hubble parameter $H$ (dotted curve). Right panel: The qualitative behavior of the $\rho_{\phi}$ (solid curve), $p_{\phi}$ (dashed curve), and $p_{{}_{\rm nc}}$ (dotted curve). Some quantities have been re-scaled for clarity. Moreover, we have assumed ${\cal W}=4$, $a_{i}=2.86$, $C=2$, $n=-1$, $\Theta=0.8$, $\Gamma=1$, $18\pi G=$ 1, and $D=4$.
Figure 4: The time behavior of the quantity $\Delta$ (see equation (22)) for $\Theta=0.0008$. We have taken the values of the other parameters as used in figure 3.
5 Commutative quantum cosmology
In this and the following sections, all variables will be treated as operators, so those with the index $c$ will be associated to the commutative quantum model, while those without an index will be related to the NC one. For the sake of simplicity, however, we omit the hats on the operators.
First, let us investigate the commutative quantum SB-FLRW cosmological model in $D$ dimensions. To get the commutative WDW equation, we should consider the canonical quantization of the classical super–Hamilton constraint (15) based on the ordinary Heisenberg-Weyl algebra:
$\left[{a}_{c},{\phi}_{c}\right]=0=\left[\Pi_{a_{c}},\Pi_{\phi_{c}}\right],$
$\left[{a}_{c},\Pi_{a_{c}}\right]=i=\left[{\phi}_{c},\Pi_{\phi_{c}}\right]~{},$ (47)
where $\Pi_{a_{c}}$ and $\Pi_{\phi_{c}}$ are the fundamental momentum operators conjugate to ${a_{c}}$ and ${\phi_{c}}$, respectively.
The canonical quantization of the Hamiltonian constraint (15) gives the WDW equation for the wave function of the universe associated with the commutative case as
$\left[-\Pi_{a_{c}}^{2}+(D-1)(D-2)\mathcal{W}^{-1}a_{c}^{-2}\phi_{c}^{-n}\Pi_{\phi_{c}}^{2}\right]\psi(a_{c},\phi_{c})$
$+\left[4(D-1)(D-2)a_{c}^{2(D-2)}V(\phi_{c})\right]\psi(a_{c},\phi_{c})=0,$ (48)
where we assumed there is no ordinary matter.
By substituting $\Pi_{a_{c}}$ and $\Pi_{\phi_{c}}$ from
$\Pi_{a_{c}}^{2}\rightarrow-a_{c}^{-p}\frac{\partial}{\partial a_{c}}\left(a_{c}^{p}\frac{\partial}{\partial a_{c}}\right),\hskip 22.76219pt\Pi_{\phi_{c}}^{2}\rightarrow-\frac{{\partial}^{2}}{\partial\phi_{c}^{2}},$ (49)
into (48) we get
$\left[\frac{\partial^{2}}{\partial a_{c}^{2}}+\frac{p}{a_{c}}\frac{\partial}{\partial a_{c}}-\frac{(D-1)(D-2)}{\mathcal{W}a_{c}^{2}}\phi_{c}^{-n}\frac{\partial^{2}}{\partial\phi_{c}^{2}}\right]\psi(a_{c},\phi_{c})\\ +\left[4(D-1)(D-2)\mathcal{V}_{0}^{2}a_{c}^{2D-4}V(\phi_{c})\right]\psi(a_{c},\phi_{c})=0,$ (50)
where $p$ is an ordering parameter that is included for the factor-ordering ambiguity [36]. In Ref. [37], assuming $p=-1$, a quantum creation of universes with compact spacelike sections corresponding to a particular case of the model (50), where $D=4$, $n=0$ and $\mathcal{W}=1/2$ has been studied in detail.
Throughout this work we will apply the simplest factor ordering of the operators, i.e. we set $p=0$.
In this section, let us proceed our investigation with an exponential scalar potential. For simplicity, we use a suitable transformation as
${a_{c}}=e^{\alpha},$ (51)
(where $\alpha=\alpha_{c}$ is a commutative variable) hence the corresponding conjugate momenta are related as
$\Pi_{a_{c}}=e^{-\alpha}\Pi_{\alpha}.$ (52)
Therefore, the WDW equation is rewritten as
-- $\left[\frac{\partial^{2}}{\partial\alpha^{2}}-\frac{(D-1)(D-2)}{\mathcal{W}}\phi_{c}^{-n}\frac{\partial^{2}}{\partial\phi_{c}^{2}}\right]\psi(\alpha,\phi_{c})$ (53)
$+$ $4(D-1)(D-2)\exp[2(D-1)\alpha]V(\phi_{c})\psi(\alpha,\phi_{c})=0~{}.$
Let us restrict ourselves to the case $n\neq-2$ and use the same exponential potential obtained in Section 3, i.e., equation (38), where for simplicity we set $\phi_{i}=0$. Using another transformation
$\phi_{c}\equiv\left[\frac{(D-1)(D-2)}{\mathcal{W}}\right]^{\frac{1}{n+2}}\varphi,$ (54)
(where $\varphi=\varphi_{c}$ is a commutative variable) the WDW equation is written as
$\left[-\frac{\partial^{2}}{\partial\alpha^{2}}+\varphi^{-n}\frac{\partial^{2}}{\partial\varphi^{2}}+4(D-1)(D-2)e^{2(D-1)\alpha}V(\varphi)\right]\psi(\alpha,\varphi)=0~{},$ (55)
where
[…]
Employing transformations
$x=-2(D-1)\alpha+\sigma\varphi^{\frac{n+2}{2}},\hskip 28.45274pty=-\alpha+\frac{8(D-1)}{\sigma(n+2)^{2}}\varphi^{\frac{n+2}{2}},$ (57)
it is easy to show that equation (53) reduces to
$\left[\frac{\partial^{2}}{\partial x^{2}}-\frac{\partial^{2}}{\partial\tilde{y}^{2}}-\upsilon e^{-x}\right]\psi(x,\tilde{y})=0~{},$ (58)
where
[…]
Assuming $\psi=X(x)Y(\tilde{y})$, equation (58) yields
$\frac{1}{X}\frac{d^{2}X}{dx^{2}}-\upsilon e^{-x}=\frac{1}{Y}\frac{d^{2}Y}{d\tilde{y}^{2}}$ (60)
The solutions for these differential equations are:
$X(x)=\begin{cases}J_{i\eta}\left(\pm 2\sqrt{\upsilon}e^{-\frac{x}{2}}\right)+J_{-i\eta}\left(\pm 2\sqrt{\upsilon}e^{-\frac{x}{2}}\right),\hskip 14.22636pt{\rm for}\hskip 14.22636pt|(n+2)\sigma|<2(D-1),\\ \\ I_{i\eta}\left(\pm 2\sqrt{\upsilon}e^{-\frac{x}{2}}\right)+K_{-i\eta}\left(\pm 2\sqrt{\upsilon}e^{-\frac{x}{2}}\right),\hskip 14.22636pt{\rm Otherwise},\end{cases}$ (61)
$Y(\tilde{y})=C_{1}\cos\left({\frac{\eta\tilde{y}}{2}}\right)+C_{2}\sin\left({\frac{\eta\tilde{y}}{2}}\right),$ (62)
where $C_{1}$ and $C_{2}$ are integration constants, $\eta$ is a separation constant, $I_{i\eta}$ and $K_{i\eta}$ are the modified Bessel Functions. In order to obtain a normalizable wave function we should establish wave packets to construct a Gaussian state.
In order to obtain the classical limit of the wavefunction, let us use the WKB–like method. Let us consider
$\psi(x,\tilde{y})=\exp{\left[-S(x,\tilde{y})\right]}$ (63)
and use the standard conditions
$\left(\frac{\partial S}{\partial x}\right)^{2}>>\Big{|}\frac{\partial^{2}S}{\partial x^{2}}\Big{|},\hskip 28.45274pt\left(\frac{\partial S}{\partial\tilde{y}}\right)^{2}>>\Big{|}\frac{\partial^{2}S}{\partial\tilde{y}^{2}}\Big{|}.$ (64)
Therefore, equation (58) reduces to
$\left(\frac{\partial S}{\partial x}\right)^{2}-\left(\frac{\partial S}{\partial\tilde{y}}\right)^{2}-\upsilon e^{-x}=0.$ (65)
Assuming $S=S_{x}(x)S_{\tilde{y}}(\tilde{y})$, the solutions of (65) are:
$S_{\tilde{y}}=s={\rm constant},\hskip 28.45274ptS_{x}=\pm\frac{2}{s\sqrt{\upsilon}}e^{-\frac{x}{2}}$ (66)
and therefore
$\Pi_{\tilde{y}}\equiv\frac{\partial S}{\partial\tilde{y}}=0,\hskip 28.45274pt\Pi_{x}\equiv\frac{\partial S}{\partial x}=\pm\frac{1}{\sqrt{\upsilon}}e^{-\frac{x}{2}}.$ (67)
Reconsidering the variables $\alpha$ and $\varphi$, we can easily obtain
$\Pi_{\alpha}=\pm\frac{2(D-1)}{\sqrt{\upsilon}}\exp{\left[(D-1)\alpha-\frac{\sigma}{2}\varphi^{\frac{n+2}{2}}\right]},$ (68)
$\Pi_{\varphi}=\mp\left[\frac{\sigma(n+2)}{2}\right]\varphi^{\frac{n}{2}}\exp{\left[(D-1)\alpha-\frac{\sigma}{2}\varphi^{\frac{n+2}{2}}\right]}.$ (69)
Finally, using equations (5), (51), (52) and (54), after some manipulation, we get back the same results for scale factor and scalar field obtained in section 3, see equations (33) and (36) for the case $n\neq-2$ with $\phi_{i}=0$.
6 Noncommutative quantum cosmology
As mentioned, solving the NC-WDW equation associated with the model investigated in Section 2 is a very complicated procedure. Therefore, from the beginning, let us restrict our attention to a NC quantum model corresponding to $\{\Pi_{a},\Pi_{\phi}\}=\theta$, $n=0$ and $V=V_{0}={\rm constant}$. Therefore, the equations (11) and (LABEL:diff.eq4) yield a constant of motion:
${\Pi}_{\phi}+\theta a=C,$ (70)
which will assist us to proceed our discussions in this Section.
Performing the quantization of the classical algebra (8) for this particular case yields a NC Heisenburg-Weyl algebra as
$\left[{a},\Pi_{a}\right]=i,\hskip 5.69046pt\left[{\phi},\Pi_{\phi}\right]=i\hskip 5.69046pt,\hskip 5.69046pt\left[\Pi_{a},\Pi_{\phi}\right]=i\theta,$ (71)
where the other commutators vanish. We should remind that the operators without index c are those associated with the NC quantum model.
The best way to represent the algebra (71) is to use the transformations
${a}$ $=$ ${a}_{c},\hskip 99.58464pt{\phi}={\phi}_{c}~{},$ (72)
$\Pi_{a}$ $=$ $\Pi_{a_{c}}+\frac{\theta}{2}{\phi}_{c},\hskip 56.9055pt\Pi_{\phi}=\Pi_{\phi_{c}}-\frac{\theta}{2}{a}_{c}~{},$ (73)
which restores the standard Heisenberg algebra. Employing transformations (72) and (73) we may consider (71) as an algebra of operators that act on the usual Hilbert space $L^{2}(\mathbb{R}^{2})$.
Using the transformations (73) provides a representation that we can write the NC-WDW equation as
$\left[-\left(-i\frac{\partial}{\partial{a_{c}}}+\frac{\theta}{2}\phi_{c}\right)^{2}+\frac{(D-1)(D-2)}{\mathcal{W}}a_{c}^{-2}\left(-i\frac{\partial}{\partial{\phi_{c}}}-\frac{\theta}{2}a_{c}\right)^{2}\right]\psi(a_{c},\phi_{c})$
$+4(D-1)(D-2)V_{0}a_{c}^{2(D-2)}\psi(a_{c},\phi_{c})=0~{},$ (74)
which explicitly depends on the NC parameter. To solve this complex equation, we invoke the NC quantum version of the constant of motion (70), which can be also written as
${C}={\Pi}_{\phi_{c}}+\frac{\theta}{2}{a}_{c},$ (75)
where we used the transformations (73). It is straightforward to show that ${C}$ commutes with NC Hamiltonian (in the constraint space of states):
$[{\Pi_{\phi}}+\theta{a},H]=0.$ (76)
Namely, any solution of equation (74) is also a simultaneous eigenstate of $C$. Let us consider $\psi_{m}(a_{c},\phi_{c})$ as an eigenstate of the operator ${C}$ with eigenvalue $m\in\mathbb{R}$:
$\left(-i\frac{\partial}{\partial\phi_{c}}+\frac{\theta}{2}a_{c}\right)\psi_{m}(a_{c},\phi_{c})=m\psi_{m}(a_{c},\phi_{c})~{},$ (77)
where we used (75).
It is easy to show that a solution for (77) is:
$\psi_{m}(a_{c},\phi_{c})=A(a_{c})\exp{\left[i\left(m-\frac{\theta}{2}a_{c}\right)\phi_{c}\right]}~{}.$ (78)
Substituting $\psi_{m}(a_{c},\phi_{c})$ from (78) into (74) yields a second order ordinary differential equation:
$A^{\prime\prime}(a_{c})+(D-1)(D-2)\left[\mathcal{W}^{-1}a_{c}^{-2}\left(m-\theta a_{c}\right)^{2}+4V_{0}a^{2(D-2)}\right]A(a_{c})=0~{},$ (79)
where a prime denotes a differentiation with respect to its argument. It is seen that equation (79), in turn, depends on the NC parameter $\theta$ and the eigenvalue $m$.
Figure 5: Numerical solution of Eq. 79, with initial conditions $A(1)=0.1$, $A^{\prime}(1)=0.1$. Moreover, we have assumed $m=1$, $\theta=-0.01$, $V_{0}=1$, ${\cal W}=2$ and $D=4$.
The equation (79) cannot be solved analytically except under certain circumstances. Let us first investigate such particular analytical solutions. In the limit of very small $a_{c}$, keeping the terms of order up to $\mathcal{O}(a_{c}^{-2})$, Eq. (79) reduces to
$A^{\prime\prime}(a_{c})+\left[(D-1)(D-2)\mathcal{W}^{-1}a_{c}^{-2}m^{2}\right]A(a_{c})$ $=$ $0~{},\hskip 14.22636pt{\rm for}\hskip 14.22636ptm\neq 0,$ (80)
$A^{\prime\prime}(a_{c})+\left[(D-1)(D-2)\mathcal{W}^{-1}\theta^{2}\right]A(a_{c})$ $=$ $0\displaystyle 0~{},\hskip 14.22636pt{\rm for}\hskip 14.22636ptm=$ 0 . (81)
For the latter, we can see that the full wave function $\psi_{m}(a_{c},\phi_{c})$ will carry effects from the NC parameter.
The analytical solution for Eq. (80) is:
$A(a_{c})=a_{c}^{\frac{1}{2}-\frac{m\sqrt{(D-1)(D-2)}\sqrt{\frac{\mathcal{W}}{m^{2}(D-1)(D-2)}-4}}{2\sqrt{\mathcal{W}}}}\left(c_{2}\>a_{c}^{\frac{m\sqrt{(D-1)(D-2)}\sqrt{\frac{\mathcal{W}}{m^{2}(D-1)(D-2)}-4}}{\sqrt{\mathcal{W}}}}+c_{1}\right),$ (82)
where $c_{1}$ and $c_{2}$ are integration constants, and can be adjusted to make $\psi_{m}(a_{c},\phi_{c})\rightarrow 0$ for $a_{c}\rightarrow 0$. Moreover, the analytical solution for Eq. (81) is given by
$A(a_{c})=c_{2}\sin\left(\sqrt{3}\theta a_{c}\right)+c_{1}\cos\left(\sqrt{3}\theta a_{c}\right).$ (83)
In the left panel of Figure 6, we depict some particular solutions for Eqs. (82) and (83) to cover the values for the eigenvalue $m=0,1,\ldots,5$. We adjusted the integration constants $c_{1}$ and $c_{2}$ such that the wave function starts at zero for $a_{c}=0$. We have, initially, an oscillatory behaviour that gets amplified as $a_{c}$ grows for all the modes (different $m$).
Furthermore, for very large $a_{c}$, by keeping only the term of order $\mathcal{O}(2(D-2))$, Eq. 79 reduces to
$A^{\prime\prime}(a_{c})+\left[4(D-1)(D-2)V_{0}a^{2(D-2)}\right]A(a_{c})=0~{}.$ (84)
An analytical solution for (84) is:
$\begin{split}A(a_{c})=\frac{\sqrt[12]{(D-1)(D-2)}\sqrt{a_{c}}}{\sqrt[6]{3}}\Biggl{[}&c_{1}\Gamma\left(\frac{5}{6}\right)J_{-\frac{1}{6}}\left(\frac{2}{3}\sqrt{(D-1)(D-2)}a_{c}^{3}\right)\\ +&c_{2}\Gamma\left(\frac{7}{6}\right)J_{\frac{1}{6}}\left(\frac{2}{3}\sqrt{(D-1)(D-2)}a_{c}^{3}\right)\Biggr{]}.\end{split}$ (85)
We observe that this solution does not depend on either the NC parameter $\theta$ or the eigenvalue $m$ for large values of $a_{c}$. Moreover, according to the right panel of Figure 6, where a solution of Eq. 85 was depicted, the behavior of the analytical solutions for large $a_{c}$ do match with the behaviour found for the numerical solution of Eq. 79 shown in Figure 5.
Figure 6: In the left panel we represent analytical solutions of Eqs. (82)-(83) for the eigenvalue $m=0,1,\ldots,5$. In the right panel we depict the behaviour of the analytical solution 85. Moreover, we have also assumed $\theta=-0.01$, $V_{0}=1$, ${\cal W}=2$ and $D=4$.
Our numerical approach investigations have shown that the solutions of equation (79) is very sensitive to the initial conditions and model parameters, as expected. Meanwhile, it is worth noting that the choice of the eigenvalue $m$ is dependent on the other CIs, thus we cannot consider the same value for it in the commutative and NC models except in one particular case. More specifically, the eigenvalue $m$ must be determined from Equation (70) assuming the classical ICs for ${\Pi}_{\phi}(0)$ and $a(0)$ which are employed to generate the solutions for equations (11)-(LABEL:diff.eq4). The only scenario in which the eigenvalue $m$ is the same for both commutative and NC instances is when $a(0)=0$. However, ignoring this particular case, we always get different values for $m$ in the commutative and NC models. For example, assuming $\Pi_{phi}(0)=10$, $a(0)=18$, and $\theta=-0.5$, we get $m=10$ for the commutative case and $m=1$ for the NC case. Figure 7 depicted this case. It is worth noting that adopting an NC scenario with an NC parameter of order one is consistent with the fundamental quantum gravity scale.
Figure 7: The he behavior of the $A(a_{c})$ (see Eq. (79)) for the commutative case (the left panel) and NC model (the right panel). We have assumed $\theta=-0.5$, $V_{0}=0.4$, ${\cal W}=2$, $D=4$, $\Pi(0)=10$, $a(0)=18$, $A(0.1)=5=A^{\prime}(0.1)$.
In the next section, we present some comments regarding our herein NC quantum case.
7 Summary and discussions
In this paper, we considered the FLRW metric and an extended version of the standard SB theory as background geometry and underlying framework in arbitrary dimensions. By introducing a dynamical deformation in the phase space and using the Hamiltonian formalism, the NC field equations were obtained, which reduce to their corresponding commutative counterparts when the NC parameter vanishes.
Our main goal was to study the evolution of the universe at early times in a (semi)classical and quantum regimes based on commutative and NC models. In this regard, we have tried to simplify our model while maintaining its realism. For example, we have just considered the spatially flat FLRW metric and assumed that the scalar field is dominant. Furthermore, we hypothesized that the scalar potential has a specific form that is useful in the classical case.
Regarding the classical NC case, using the scalar potential (25) (that leads to the power-law inflation in the standard models), we have applied two methods to analyze the behavior of the physical quantities. In the first method, we used the numerical analysis, which gives the following results for the small values of the NC parameter. (i) In the earlier times, the scale factor is accelerated, so after a very short time it is decelerated. The later epoch can be assigned to the universe dominated by radiation. This successful graceful exit obtained by the presence of NC effects. (ii) With regard to the horizon problem, we have considered the nominal condition, which is satisfied for our classic NC model. These results indicate that our NC scenario can be considered as a successful inflationary model. (iii) At late times, $\ddot{a}$ asymptotically approaches zero, which can be attributed to the coarse-grained explanation by quantum gravity effects. In Section 4, by constructing a suitable configuration for the dynamical system, we plotted the phase portrait of equations (40) and (41) that confirms the above consequences. (iv) At the level of the field equations, we obtained an evolutionary equation for the scale factor to compare it with the Starobinsky model. It seems that the NC parameter can play a crucial role in finding a correspondence between these models.
However regarding the second method of our herein NC classical model, we took the ansatz (44) as the solution of the field equations that leads to an exact solution for the scalar field. Our numerical investigations demonstrate that considering the solutions for the scale factor and the scalar field associated to this method indicates that the modified Klein–Gordon equation is still valid (or equivalently the conservation of the EMT (22) is satisfied) only when the NC parameter takes very small values. In what follows, let us mention some interesting features of these solutions. (i) The evolution of the scale factor bears close resemblance with that obtained in pre-Big-Bang scenarios [35]: concretely, from (44), we obtain $\dot{H}(t)=-\frac{t}{\Theta}H(t)$. (ii) We can easily show that the number of e-folds is given by $\mathcal{N}=H_{i}\sqrt{2\pi\Theta},$ as where we assumed the asymptotic values of the scale factors $a_{i}$ and $a_{f}$ at $t=\pm\infty$. According to the corresponding field equations, in four dimensions, we have $H_{i}=(8\pi G\rho_{i}/3)^{1/2}$, where $\rho_{i}$ is the value of the energy density of the universe at early times. For instance, assuming $\mathcal{N}\sim 60$, we obtain $\sqrt{\Theta}\sim 3.6l_{pl}$, where $l_{pl}$ denotes the Planck length. This results indicates that when only the scalar field dominates, the number of e-folds depends on the non-commutative parameter which is of order the Planck length.
In Section 5, we have investigated the commutative quantum cosmology in arbitrary dimensions. We have obtained solutions for the WDW equation analytically for the same potential as in the classical models. The classical limit of the wave function was then found using the WKB approximation, and it was demonstrated that these solutions are the same as those obtained in Section 3.
At the NC quantum level, we focused our research on a specific but fascinating situation. In contrast to the extremely complicated form of the NC-WDW equation, we have shown that there is a constant of motion that commutes with the Hamiltonian in the constraint space of states. The strategy of exploiting this important conclusion helped transform the NC–WDW equation into an ordinary differential equation that can be solved numerically in the general case or analytically in approximate cases.
In the general scenario, eigenvalue $m$ cannot have the same value for the commutative and NC models. It is obtained in more concrete terms by choosing the classical initial values for the scale factor and momentum associated with the scalar field. Our numerical research have revealed that for the NC scenario, damping behavior in the wave function is always conceivable for values of the NC parameter close to one. Because of the postulated noncommutativity of the momenta, such intriguing quantum characteristics of the very early universe were obtained.
It is important to note that the noncommutativity we have chosen for this work (in both the classical and quantum regimes) is not unique. For instance, in our model, in addition to the noncommutativity between momenta, we can also propose a noncommutativity between the scale factor and the scalar field. When this happens, the equations undoubtedly become more complex, leading to greater degrees of freedom and ultimately more appropriate solutions. We should mention that the amplitude $|\Psi(a,\phi)|^{2}$ may be replaced with another quasiprobability distribution [26] that is generated using the Moyal product in order to obtain probability measures in the NC models.
Finally, it is worth noting that the NC cosmological models have both advantages and shortcomings (some of which will be summarized in the following paragraphs, respectively), compared to the corresponding commutative models in the classical and quantum regimes.
NC models have been used to eliminate singularities such as the big bang, which were anticipated by the corresponding commutative cosmological model. These models can lead to an inflationary phase in the early universe (without using ad hoc inflationary potentials, as in standard models) and the quantum gravity effects more effectively than the commutative models, resulting in a more complete description of the universe at small scales. They may also produce novel observables and provide opportunities for observational verification.
The field equations associated with the NC models are more complicated to solve than those of the corresponding commutative models, making mathematical analysis and calculations more difficult. Moreover, physical interpretation of the consequences related to the NC models is challenging because they may add atypical aspects that are difficult to understand. Furthermore, NC models have not been experimentally tested, making it difficult to determine their validity and compare them to the standard models.
Acknowledgments
SMMR acknowledges the FCT grants UID-B-MAT/00212/2020 and UID-P-MAT/00212/2020 at CMA-UBI plus the COST Action CA18108 (Quantum gravity phenomenology in the multi-messenger approach).